OpenStudy (anonymous):

Choose one of the factors of 1331x^3 - 8y^3. 4 11x - 2y 121x2 + 44xy + 4y2 121x2 - 22xy + 4y2 Question 4 Choose one of the factors of x^3 + 343. x - 7 x2 - 14x + 49 x2 + 7x + 49 None of the above

OpenStudy (anonymous):

@e.mccormick

OpenStudy (anonymous):

@Preetha

OpenStudy (anonymous):

help me plz

jhonyy9 (jhonyy9):

so you get that 11*11*11=1331 and 2*2*2=8 this mean that 1331x^3 -8y^3 is a different of perfect cubes do you know formula for this so a^3 -b^3 = ?